# ASiM: Improving Transparency of SRAM-based Analog Compute-in-Memory Research with an Open-Source Simulation Framework

### v2

@@ III-B6 ADC Modeling
$MAC=\text{clamp}\left(\left\lfloor V_{MAC}\right\rceil,0,2^{k}-1\right)$ (5)
[…]
@@
III-C Noise-Aware Training
@@ IV Performance Analysis of ACiM Inference
IV-A Re-consider CSNR with Actual Data Distribution
@@ IV-B ADC Quantization Noise
We begin our exploration of inference accuracy by examining the quantization noise introduced by ADCs in ACiM circuits. Unlike DCiM, where inference accuracy is influenced solely by model quantization, the ADC rounding effect in ACiM circuits introduces additional quantization errors to each MAC output.
@@ IV-C Analog Noise
Although our previous analysis demonstrated that ADC quantization error minimally impacts DNN inference accuracy, making bit-parallel scheme viable for ACiM efficiency improvements, this assumption is established on ideal noise-free conditions. Real ACiM accelerators face various additional noise sources that require careful consideration. To assess this, we incorporated random noise and non-linearity into simulations, analyzing their effects on inference accuracy. As detailed in Section III, random noise includes $kT/C$ noise and comparator decision errors, while non-linearity accounts for layout-specific issues and device mismatch, analyzed here based on capacitor mismatch in the CBL and DAC.
@@
In the non-linearity analysis shown in Fig. 9(b), a similar trend appears, with the transition boundary shifting slightly downward as non-linearity diminishes with increased ADC output code. However, as seen in Fig. 5, the MAC output is concentrated in lower-valued regions, making non-linearity effects still significant.
Figure 9: Shmoo plot of inference accuracy across different noise intensities for an ACiM macro with a row-parallelism of 256 using an 8-bit ADC, evaluated by an 8b/8b DNN model. Two noise models are assessed separately: (a) Random noise and (b) Non-linearity. The accuracy of the digital counterpart is provided alongside each model’s name for reference.
@@ IV-D Simulation to Silicon Correlation
Silicon Validation. To validate the accuracy of ASiM in predicting inference results for real chips, we compared its ResNet-18 accuracy predictions with the silicon validation results of our prototype chip reported in CR-CIM [35]. In the silicon test, CR-CIM processed the compute-intensive CONV and Linear layers of an 8b/8b ResNet-18 on-chip, while the remaining layers were handled off-chip, replicating the same setup used in ASiM simulations. The measured $\text{LSB}_{\text{rms}}$ noise was equally divided and allocated into random noise and non-linearity. ASiM predicted an accuracy of 91.84%, which closely corresponded to the 91.70% measured in CR-CIM, verifying ASiM’s correctness in evaluating inference accuracy.
@@ V-A Hybrid CiM
ACiM begins to outperform DCiM at a row-parallelism of 64, with energy efficiency improving to a 7×7$\times$7 × gain as row-parallelism scales up to 1024 [4]. At a row-parallelism of 256, ACiM achieves approximately 3×3$\times$3 × the TOPS/W of DCiM. Building upon this benchmark, the HCiM approach illustrated in Fig. 13(b) results in 6%, 20%, 30%, and 36% decrease in TOPS/W for CIFAR-10/ResNet-18, ImageNet/ResNet-18, CIFAR-10/ViT-B-32, and ImageNet/ViT-B-32, respectively, compared to full ACiM execution. However, HCiM still provides 182%, 140%, 110%, and 90% TOPS/W gains over full DCiM execution, showcasing its strong potential for energy-efficient CiM deployment in real-world applications.
TABLE I: Evaluation of Different Simulation Frameworks for DNN Inference Assessment.
@@
Aligned with Realistic ADC $\times$ $\times$ $\times$ $\times$ $\checkmark$
@@ VI-A Comparison with Other Simulation Frameworks
Table I compares our ASiM framework with other inference simulation frameworks. One key distinction lies in the modeling of ADC behavior, which significantly impacts inference performance. Saikia’s modeling [22] employs an adaptive dynamic range, which is challenging to implement in real-world applications, whereas CIMUFAS [21] bypasses ADC rounding by introducing Gaussian noise instead. NeuroSim [14, 39] assumes that the ADC dynamic range is defined by the Min-Max range of each MAC output, leading to overly optimistic inference results, when real data distributions such as Fig. 5 are considered. Unlike these frameworks, ASiM simulates ADC behavior across its full dynamic range, accurately capturing both rounding effects and noise-induced errors based on the real MAC output distribution. Additionally, ASiM supports the latest Transformer-based models and incorporates advanced features such as bit-parallel ACiM evaluation. Lastly, ASiM’s modular design allows for seamless integration into various codebases, making it a user-friendly and adaptable tool for ACiM circuit evaluation.
@@ VI-B Model Extension
As ACiM technology rapidly evolves, novel designs continue to enhance area efficiency, energy efficiency, and throughput. Although ASiM is primarily developed for charge-domain SRAM ACiM, it serves as a flexible framework that users can modify to model emerging architectures. For example, adjusting the execution order in Algorithm 1 (moving lines 7–10 after line 13) enables support for weight bit-parallel computations [10, 6]. In current domain SRAM ACiM, replacing capacitance mismatch based non-linearity model with gain compression can effectively capture the non-linearity introduced by current roll-off when transistors transition from saturation to the triode region. Further adding random variations to binary weight tensors enables accurate modeling of local transistor variation in current-domain computations. We encourage users to extend ASiM to reflect their novel circuit designs, ensuring alignment with their specific hardware characteristics.
@@ VII Conclusions
We presented ASiM in this paper, an open-source simulation framework specifically developed to assess and optimize inference accuracy in SRAM-based ACiM circuits. ASiM allows researchers to easily integrate their ACiM circuits into the PyTorch ecosystem, enabling the evaluation of various design factors under realistic conditions that include quantization noise, analog imperfections, and ADC constraints.
@@
Our extensive analysis using ASiM revealed critical insights for ACiM design. Although the bit-parallel approach can substantially improve energy efficiency while tolerating quantization noise, analog noise poses a serious threat to inference accuracy, creating stringent design requirements. We also showed that higher ADC precision is crucial for more complex models and tasks, such as ImageNet and ViT, while CNNs on simpler datasets can reduce ADC precision without significant accuracy loss. Furthermore, we found that largest computation errors primarily occur in the MSB cycles, where even a minor ADC readout error as small as 1 LSB can result in severe inference degradation.
To tackle these challenges, we recommended two strategies: HCiM and majority voting, both of which aim to ensure accurate MSB cycle computation. Incorporating these methods shows promise in enhancing inference accuracy without diminishing inherent advantage on energy efficiency, presenting a promising strategy for deploying ACiM reliably in real-world applications.

### v3

@@ III-B6 ADC Modeling
$MAC=\text{clamp}\left(\left\lfloor V_{MAC}\right\rceil,0,2^{k}-1\right),$ (5)
where $V_{MAC}$ denotes the CBL voltage before rounding, and $MAC$ represents the digital MAC value after ADC rounding. Here, $k$ is the ADC precision, and $\left\lfloor\cdot\right\rceil$ denotes the round-to-nearest operator. After rounding, the digital output is scaled back to its corresponding MAC value, followed by bit-shifting and accumulation to produce the final multi-bit output activations.
@@
TABLE I: Comparison of inference accuracy for a bit-serial ACiM macro with 256-row parallelism under a 6-bit ADC and 0.5 $\text{LSB}_{\text{rms}}$ noise, evaluated using several existing simulators with an 8b/8b ResNet-18 model on ImageNet.
CSNR CIMUFAS Saikia NeuroSim
ADC Model SNR Gaussian Adaptive Min-Max
Accuracy 65.44% 0.10% 40.82% 61.09%
III-C Preliminary Analysis of Existing Simulators
We conducted a preliminary analysis using several prevailing simulation frameworks, and the predicted results are summarized in Table I. While each framework has its merits, differences in modeling approaches lead to divergent accuracy predictions, underscoring the importance of SRAM-specific modeling capabilities as offered by ASiM.
Our analysis identifies ADC modeling and noise representation as key factors driving the divergence in accuracy predictions. For instance, the CSNR approach [24, 30, 10] models quantization and analog noise by injecting scaled random noise into the model’s forward pass based on a predefined SNR. However, it neglects the noise amplification caused by shifting and accumulation of partial MAC results, resulting in significant discrepancies from actual SRAM-based ACiM circuit behavior. While CIMUFAS [20] performs binary-level decomposition of CiM operations, it does not include an explicit ADC model and instead introduces Gaussian noise as a substitute, potentially limiting its ability to accurately reflect inference performance. Saikia’s model [29] assumes an idealized adaptive dynamic range mechanism that is not readily implementable in real ADCs, thereby underestimating the quantization error introduced by degraded ENOB in practical hardware. NeuroSim [14, 31] quantizes ADC outputs based on the Min-Max range of MAC results, which leads to overly optimistic estimations of ADC rounding effects and the corresponding impact of analog noise. Although each framework has its strengths, these modeling differences can lead to diverging accuracy predictions. Without clear guidance, such discrepancies may inadvertently influence circuit design decisions. This highlights the need for a consistent and accurate simulation methodology to support reliable evaluation and design exploration for ACiM systems.
To address these concerns, ASiM simulates ADC behavior across its full dynamic range, accurately capturing both rounding effects and noise-induced errors based on the real MAC output distribution. Additionally, ASiM supports the latest Transformer-based models and incorporates advanced features such as bit-parallel ACiM evaluation. Lastly, ASiM’s modular design allows for seamless integration into various codebases, making it a user-friendly and adaptable tool for ACiM circuit evaluation.
III-D Noise-Aware Training
@@ IV Performance Analysis of ACiM Inference
IV-A Re-evaluating CSNR Under Actual Data Distributions
@@ IV-B ADC Quantization Noise
We begin our exploration of inference accuracy by examining the quantization noise introduced by ADCs in ACiM circuits. Unlike Digital Compute-in-Memory (DCiM), where inference accuracy is influenced solely by model quantization, the ADC rounding effect in ACiM circuits introduces additional quantization errors to each MAC output.
@@ IV-C Analog Noise
Although our previous analysis demonstrated that ADC quantization error minimally impacts DNN inference accuracy, making bit-parallel scheme viable for ACiM efficiency improvements, this assumption is established on ideal noise-free conditions. Real ACiM accelerators face various additional noise sources that require careful consideration. To assess this, we incorporated random noise and nonlinearity into simulations, analyzing their effects on inference accuracy. As detailed in Section III, random noise includes $kT/C$ noise and comparator decision errors, while nonlinearity accounts for layout-specific issues and device mismatch, analyzed here based on capacitor mismatch in the CBL and DAC.
@@
In the nonlinearity analysis shown in Fig. 9(b), a similar trend appears, with the transition boundary shifting slightly downward as nonlinearity diminishes with increased ADC output code. However, as seen in Fig. 5, the MAC output is concentrated in lower-valued regions, making nonlinearity effects still significant.
Figure 9: Shmoo plot of inference accuracy across different noise intensities for an ACiM macro with a row-parallelism of 256 using an 8-bit ADC, evaluated by an 8b/8b DNN model. ResNet-18 is trained with 100% and 50% , and ViT-B-32 is trained with 60% and 40% training noise intensities for CIFAR-10 and ImageNet, respectively. Two noise models are assessed separately: (a) Random noise and (b) Nonlinearity. The accuracy of the digital counterpart is provided alongside each model’s name for reference.
@@ IV-D Simulation to Silicon Correlation
Silicon Validation. To validate the accuracy of ASiM in predicting inference results for real chips, we compared its ResNet-18 accuracy predictions with the silicon validation results of our prototype chip reported in CR-CIM [30]. In the silicon test, CR-CIM processed the compute-intensive CONV and Linear layers of an 8b/8b ResNet-18 on-chip, while the remaining layers were handled off-chip, replicating the same setup used in ASiM simulations. The measured $\text{LSB}_{\text{rms}}$ noise was equally divided and allocated into random noise and nonlinearity. ASiM predicted an accuracy of 91.84%, which closely corresponded to the 91.70% measured in CR-CIM. Furthermore, we configured our prototype chip with varying ADC precision levels, and ASiM successfully identified the inflection point at which inference accuracy begins to degrade. The predicted trend closely matches our silicon measurements, supporting the validity of ASiM in evaluating inference accuracy.
@@ V-A Hybrid CiM
ACiM begins to outperform DCiM at a row-parallelism of 64, with energy efficiency improving to a $7\times$ gain as row-parallelism scales up to 1024 [4]. At a row-parallelism of 256, ACiM achieves approximately $3\times$ the TOPS/W of DCiM. Building upon this benchmark, the HCiM approach illustrated in Fig. 13(b) results in 6%, 20%, 30%, and 36% decrease in TOPS/W for CIFAR-10/ResNet-18, ImageNet/ResNet-18, CIFAR-10/ViT-B-32, and ImageNet/ViT-B-32, respectively, compared to full ACiM execution. However, HCiM still provides 182%, 140%, 110%, and 90% TOPS/W gains over full DCiM execution, showcasing its strong potential for energy-efficient CiM deployment in real-world applications.
TABLE II: Comparison of Different Simulation Frameworks for DNN Inference Assessment.
@@ VI-A Comparison with Other Simulation Frameworks
Table II compares our ASiM framework with other inference simulation frameworks. One key distinction lies in the modeling of ADC behavior, which significantly impacts inference performance. Saikia’s modeling [29] employs an adaptive dynamic range, which is challenging to implement in real-world applications, whereas CIMUFAS [20] bypasses ADC rounding by introducing Gaussian noise instead. NeuroSim [14, 31] assumes that the ADC dynamic range is defined by the Min-Max range of each MAC output, leading to overly optimistic inference results. Unlike these frameworks, ASiM simulates ADC behavior across its full dynamic range, accurately capturing both rounding effects and noise-induced errors based on the real MAC output distribution. Additionally, ASiM supports the latest Transformer-based models and incorporates advanced features such as bit-parallel ACiM evaluation. Lastly, ASiM’s modular design allows for seamless integration into various codebases, making it a user-friendly and adaptable tool for ACiM circuit evaluation.
@@ VI-B Model Extension
As ACiM technologies rapidly evolve, the incorporation of emerging memories and novel circuit designs continues to enhance system efficiency. Although ASiM is initially designed for charge-domain SRAM ACiM, its modular and adaptable structure allows users to easily modify it for other architectures. Taking RRAM as an example, where crossbar cells store multi-bit weights, the loop over binary weight bits in Algorithm 1 is unnecessary and can be omitted. This adjustment enables each ADC to translate the quantized current value for a RRAM column. Meanwhile, the capacitor nonlinearity model used for SRAM should be replaced with a noise model that more accurately captures conductance variation in RRAM. With such minimal changes, users can leverage ASiM’s core logic to create more faithful simulations for their specific circuit implementations. We encourage the research community to extend ASiM to better support diverse and emerging ACiM technologies.
VI-C Comparison across State-of-the-Art CiM designs
To evaluate the practical inference performance of SRAM-based ACiM, we compare representative designs that can be accurately modeled within the ASiM framework. All simulations are conducted using a unified 8b/8b ResNet-18 model on the CIFAR-10 benchmark. The resulting inference accuracy, energy efficiency, and area efficiency are normalized to a 22 nm technology node and 0.9 V supply voltage and summarized in Fig. 15. The size of each bubble represents the estimated area efficiency, calculated as the average area per basic CiM cell. Relevant design parameters are extracted from their respective publications and summarized in Table III. For designs where $\text{LSB}_{\text{rms}}$ was not reported, a nominal value of 0.5 LSB was assumed. The designs under evaluation include a range of representative schemes: the standard Bit-Parallel/Bit-Serial (BPBS [24]) architecture for multi-bit DNN processing; the Sparsity-Adaptive (Spar.-Adapt. [13]) scheme, which skips redundant ADC conversions by exploiting activation sparsity; the End-to-End Analog (E2E Analog [40]) pipeline, which maintains analog signal integrity throughout computation; the Differential Switched-Capacitor (Diff.-SC [41]) design, which utilizes differential compute lines to enhance dynamic range and ADC precision; and the Capacitor-Reconfigured CiM (CR-CIM [30]) scheme, which reuses cell capacitors to improve ADC resolution. In addition, CR-CIM is also configured with a reduced 6-bit ADC to demonstrate accuracy degradation, followed by recovery via majority voting. For completeness, the Hybrid CiM (HCiM [42]) and two digital CiM (DCiM) designs [43, 44] are also included as references.
Figure 15: Comparison of state-of-the-art CiM designs analyzed using ASiM. Design parameters are listed in Table III. Bubble size indicates the estimated chip area, and both energy efficiency and area are normalized to a 0.9 V supply voltage at the 22 nm technology node using spatial scaling factors.
TABLE III: Parameters of state-of-the-art ACiM designs presented in Fig. 15, with specifications extracted from their publications.
Design Parallelism Act. Enc. ADC $\textbf{LSB}_{\textbf{rms}}$
BPBS [24] 2,305 1b 8b 0.37
Spar.-Adapt. [13] 256 1b 8b N/A
E2E Analog [40] 64 4b 5b 0.27
Diff.-SC [41] 1,152 4b 10b 0.40
CR-CIM [30] 512 1b 10b 0.58
Hybrid [42] 256 2b 8b N/A
Based on Fig. 15, the Spar.-Adapt. and 8-bit CR-CIM designs yield the highest inference accuracy, primarily because they operate above the boundary ADC precision. In contrast, more aggressive designs typically demonstrate superior energy efficiency but at the cost of reduced inference accuracy. Designs with lower reported $\text{LSB}_{\text{rms}}$ values tend to better support such aggressive configurations, underscoring the importance of meticulous ACiM design. The HCiM architecture proposed in [42] achieves approximately $2\times$ energy efficiency improvement and significantly reduced area cost compared to standard DCiM, while incurring negligible accuracy loss, highlighting its promise for future exploration. Furthermore, by applying 3 times oversampling to majority voting with a boundary level of 3, the 6-bit CR-CIM design recovers accuracy from 72.74% to 90.55%, while achieving a 65% improvement in energy efficiency compared to its 8-bit configuration. This indicates its strong potential for low-power applications where moderate accuracy degradation is acceptable.
@@ VII Conclusions
We presented ASiM in this paper, an open-source simulation framework specifically developed to assess and optimize inference accuracy in SRAM-based ACiM circuits. ASiM allows researchers to easily integrate their ACiM circuits into the PyTorch ecosystem, enabling the evaluation of various design factors under realistic conditions that include quantization noise, analog imperfections, and ADC constraints. Our extensive analysis using ASiM revealed critical insights for ACiM design. Although the bit-parallel approach can substantially improve energy efficiency while tolerating quantization noise, analog noise poses a serious threat to inference accuracy, creating stringent design requirements. We also showed that higher ADC precision is crucial for more complex models and tasks, such as ImageNet and ViT, while CNNs on simpler datasets can reduce ADC precision without significant accuracy loss. Furthermore, we found that largest computation errors primarily occur in the MSB cycles, where even a minor ADC readout error as small as 1 LSB can result in severe inference degradation. To tackle these challenges, we recommended two strategies: HCiM and majority voting, both of which aim to ensure accurate MSB cycle computation. Incorporating these methods shows promise in enhancing inference accuracy without diminishing inherent advantage on energy efficiency, presenting a promising strategy for deploying ACiM reliably in real-world applications.
A General Perspective on MAC Distribution
Although we only illustrate the MAC output of convolutional layers in CNNs, we argue that the trend shown in Fig. 5 is generalizable across diverse DNN models and layers. In particular, the MAC output of each ACiM column is proportional to the bit-level sparsity of weights and activations. The expected MAC value at an ACiM column can be expressed as:
$P(DP=1)=P(w=1\cap x=1)=P(w=1)\cdot P(x=1),$ (7)
$E(MAC)=n\cdot P(DP=1).$ (8)
Here, the expected MAC output equals the product of the row parallelism $n$ and the probability that a local dot product evaluates to ”1”. This probability further decomposes into the product of the bit-level sparsity of weights and activations [35]. The derivation indicates that MAC outputs are fundamentally governed by sparsity patterns. For example, even when both weight and activation bits have a 50% probability of being ”1”, the MAC output only spans roughly 25% of the full dynamic range. Thus, sparsity inherently constrains the MAC span, and this trend persists across different depths and architectures.
To further analyze distribution characteristics, the weights and input activations in DNNs generally follow two representative types: (1) a symmetric, Gaussian-like distribution, or (2) a one-sided long-tailed distribution. After quantization into binary representations, their bit-level sparsity patterns align with those shown in Fig. 4(b). For type (1), the bit sparsity across indices fluctuates around 50%, whereas for type (2), it typically falls between 0% and 40%, depending on the bit index. Although distribution shapes may vary across models and layers, the resulting bit-level sparsity after quantization remains consistent. As illustrated in Fig. 4(b), two dominant sparsity curves correspond to these two distribution types, despite some shape differences observed in Fig.4(a). Overall, modern DNN weights and activations largely conform to one of these two categories. Consequently, according to Eq. 8, the column-wise MAC voltages observed by ADCs are confined within a limited dynamic range, which warrants careful modeling for inference accuracy.
For reference, we analyze the MAC outputs of a 256-row-parallel ACiM macro in the Query Projection of the Attention Block in ViT-B-32, with both weights and activations quantized to 6 bits. As shown in Fig. 16, the observed MAC values remain consistently below 60, which is far smaller than the theoretical maximum of 256. Remarkably, this distribution is even narrower than that observed in CNN layers (Fig. 5), thereby imposing stricter precision requirements on ACiM circuits.
Quantitative Evaluation of Noise-Aware Training
In this section, we quantitatively evaluate the effectiveness of NAT. We analyze ResNet-18 and ViT-B-32 models trained on CIFAR-10 and ImageNet with noise intensities of 100%, 60%, 50%, and 40%. These models were subsequently tested under varying inference-time noise levels and compared with their non-NAT counterparts, as shown in Fig. 17. The results clearly indicate that NAT enhances model robustness, enabling improved tolerance to higher noise levels during inference.
To strengthen our argument, we simulate the 8b/8b ResNet-18 model on CIFAR-10 using the same experimental configuration as in Fig. 9(a), except that the weights are obtained without NAT. The corresponding accuracy shmoo in Fig. 18 shows a sharper decline relative to the NAT-enhanced counterpart, thereby underscoring the necessity of NAT in improving the robustness of ACiM-based systems.
Examining Noise Modeling in ASiM
Figure 19: Linearity plots of a 4-bit ADC under different noise models: (a) Random noise. (b) Nonlinearity-induced distortion. (c) Random noise with majority voting applied via 5 times oversampling.
To provide clearer insight into the noise models adopted in ASiM, we visualize the simulated noise behavior using linearity plots. Figs. 19(a) and (b) show the MAC output characteristics of a 4-bit ADC under both the random noise model and the nonlinearity model. In both cases, the relative noise intensity is configured at $\text{LSB}_{\text{rms}}=1.0$, and the error bars indicate the standard deviation ($\sigma$) of the MAC readout.
For the random noise model, each MAC output follows a Gaussian distribution with $\sigma\approx 1.02$ LSB. Under the nonlinearity model, smaller MAC values show larger variations ($\sigma\approx 1.09$ LSB), whereas the variation decreases gradually for larger MAC values ($\sigma\approx 0.17$ LSB). This trend is consistent with capacitor mismatch effects, where fewer active capacitors induce greater variation at lower MAC levels.
The majority voting approach mitigates analog noise by statistically reducing errors through oversampling. To illustrate this effect, we apply $5\times$ oversampling (with $\text{LSB}_{\text{rms}}=1.0$) and show the corresponding linearity plot in Fig. 19(c). As observed, majority voting reduces the standard deviation from approximately 1.02 LSB to 0.44 LSB, thereby improving MAC precision and enhancing inference accuracy.